OpenStudy (anonymous):

let f(x)=(sinx)^2-(cosx) on the interval (0,2pi). where on this interval does f increase and where does it decrease? I tried deriving it, but I can't seem to get it right.

OpenStudy (anonymous):

\[f'(x)=2\sin(x)\cos(x)+\sin(x)\]

OpenStudy (anonymous):

\[2\sin(x)\cos(x)+\sin(x)=0\] \[\sin(x)(2\cos(x)+1)=0\] \[\sin(x)=0, x=0,x=\pi\] \[2\cos(x)+1=0\implies \cos(x)=-\frac{1}{2}\implies x=\frac{2\pi}{3}, x=\frac{4\pi}{3}\] etc